find the slope of the line 5x +2y =1.
Found 2 solutions by stanbon, anantha:
Answer by stanbon(75887)

find the slope of the line 5x +2y =1.
--------------
Put the equation in slope-intercept form:
2y = -5x + 1
y = (-5/2)x + (1/2)
----
slope = -5/2
=====================
Cheers,
Stan H.

Answer by anantha(86)

equation of the line=
5x+2y=1
the equation is of the form ax+by+c=0 then
the slope of the line= -(x coefficient)/(y coefficient)
step1:1st we convert given line in to ax+by+c=0 for this we subtract 1 on both sides
5x+2y-1=1-1
5x+2y-1=0
now the equation is of the form ax+by+c=0
slope= -(x coefficient)/(y coefficient)
here x coefficient=5,y coefficient=2
slope= -(5)/2